OpenStudy (darkigloo):

How do I find work? I have to find how much work is needed to accelerate 1200kg from rest to a speed of 2m/s. m= 1200kg = 11772N vi=0m/s vf=2m/s I know W=Fd

OpenStudy (roadjester):

\(\LARGE W=Fdcos\theta=madcos\theta\). You're acceleration is not necessarily 9.8. You have to find the acceleration (possibly using a kinematic equation) to find the acceleration. Unless of course your force is going downwards. In which case downward acceleration is 9.8.

OpenStudy (anonymous):

600 maybe..?

OpenStudy (roadjester):

Without the \(\LARGE FULL\) question, I can't help you much.

OpenStudy (roadjester):

@Haileymae and how did you get 600? And in what units?

OpenStudy (anonymous):

I divided 1200 by 2? Probably in N

OpenStudy (roadjester):

Work is in Joules....

OpenStudy (darkigloo):

The question is: How much work is needed to accelerate a 1200kg car from rest to a speed of 2m/s. That's all I was given..

OpenStudy (roadjester):

\(\large v_f^2=v_i^2 +2a\Delta x\) is the first thing that comes to mind although you don't know relative position... Think about this, cosine of 90 is 0 so if your force is downward and displacement is forward, there is no work. That's all I got. sorry

OpenStudy (darkigloo):

:) no problem, i appreciate it. thanks.